# RuCCoD: Towards Automated ICD Coding in Russian

### v1

@@ Abstract
This study investigates the feasibility of automating clinical coding in Russian, a language with limited biomedical resources. We present a new dataset for ICD coding, which includes diagnosis fields from electronic health records (EHRs) annotated with over 10,000 entities and more than 1,500 unique ICD codes. This dataset serves as a benchmark for several state-of-the-art models, including BERT, LLaMA with LoRA, and RAG, with additional experiments examining transfer learning across domains (from PubMed abstracts to medical diagnosis) and terminologies (from UMLS concepts to ICD codes). We then apply the best-performing model to label an in-house EHR dataset containing patient histories from 2017 to 2021. Our experiments, conducted on a carefully curated test set, demonstrate that training with the automated predicted codes leads to a significant improvement in accuracy compared to manually annotated data from physicians. We believe our findings offer valuable insights into the potential for automating clinical coding in resource-limited languages like Russian, which could enhance clinical efficiency and data accuracy in these contexts.
@@
Aleksandr Nesterov1, Andrey Sakhovskiy2, Ivan Sviridov3, Airat Valiev4 Vladimir Makharev1, Petr Anokhin1, Galina Zubkova3, Elena Tutubalina1,2,5 1 AIRI, Moscow, Russia 2 Sber AI, Moscow, Russia 3 Sber AI Lab, Moscow, Russia 4 HSE University, Moscow, Russia 5 ISP RAS Research Center for Trusted Artificial Intelligence, Moscow, Russia
@@ 1 Introduction
While ICD codes assignment is crucial for EHRs, it poses significant challenges. Human coders must navigate a wide array of medical terminology, subjective interpretations, and time pressures, all while staying updated with constantly changing classification standards Burns et al. (2012); O’Malley et al. (2005); Cheng et al. (2009). Coding errors can lead to misdiagnosis, ineffective treatment, diminished trust in the healthcare system, and negative public health outcomes. Furthermore, errors in manual coding, according to the ICD system, result in financial repercussions, accounting for 6.8% of the total payments Manchikanti (2002).
@@
Despite extensive research on ICD coding using neural networks Li and Yu (2020); Zhou et al. (2021); Yuan et al. (2022a); Baksi et al. (2024); Boyle et al. (2023); Mullenbach et al. (2018a); Cao et al. (2020); Yuan et al. (2022a); Yang et al. (2022), significant challenges persist for non-English languages. These include low inter-coder agreement, limited labeled data, variability in clinical notes, the hierarchy of ICD codes, and reliance on incomplete input data. To address these issues, we introduce a novel dataset for automatic ICD coding in Russian.
@@
Previous studies have primarily focused on English-language datasets, specifically MIMIC-III/IV Goldberger et al. (2000); Johnson et al. (2023). Russian is considered underdeveloped in the clinical domain, despite being one of the top ten languages. The Russian segment of the Unified Medical Language System (UMLS) Bodenreider (2004) comprises only 1.96% of the vocabulary and 1.62% of the source counts found in the English UMLS NIH . Recent corpora, such as RuCCoN Nesterov et al. (2022a) and NEREL-BIO Loukachevitch et al. (2023), focus on concepts within the Russian UMLS. Our study shows that UMLS may not fully meet the structured requirements of ICD coding, particularly in semantic representation, limiting the transferability of models dependent on UMLS for clinical tasks (Sec. 4).
@@
In this paper, we present RuCCoD (Russian ICD Coding Dataset), a new dataset for ICD coding in Russian [^1] We will release this dataset upon acceptance., labeled by medical professionals based on concepts from the ICD-10 CM (Clinical Modification) system (Sec. 3.1). Second, we establish a comprehensive benchmark for state-of-the-art models, including a BERT-based Devlin et al. (2019) pipeline for information extraction, a LLaMa-based Touvron et al. (2023) model with Parameter Efficient Fine-Tuning (PEFT) and with retrieval-augmented generation (RAG). Additionally, we evaluate the transfer of model performance from UMLS codes and similar datasets, such as PubMed abstracts Loukachevitch et al. (2023) and clinical notes Nesterov et al. (2022a) (Sec. 4).
@@
Finally, we perform a set of experiments on a large in-house dataset of 865k EHRs from 164k patients, divided into training and testing sets (further named RuCCoD-DP). Using the best-performing model on RuCCoD, we generated ICD codes for the training set and evaluated how performance varies when employing our automatic labels for a diagnosis prediction task, compared to those assigned by doctors during patient visits (Sec. 5). Expriments have shown a great potential of automatically labeled EHRs for training a disease diagnostic model. Specifically, pre-training on automatically assigned ICD codes gives a huge macro-averaged F1-score growth of 28% for diagnosis prediction compared to physician-assigned ICD codes. The finding indicates a great complexity of diagnosis formalization within the ICD system for doctors as well as the great perspective for AI-guided diagnosis prediction. We hope our work to provide a foundation and guidance for researchers working in low-resource clinical languages.
Figure 2: Schematic description of ICD coding (in blue) and diagnosis prediction tasks (in yellow). Diagnosis prediction uses prior EHR data and current visit details, excluding the doctor’s conclusion, which is used for ICD coding to generate AI-assigned ICD codes. Both original and AI ICD code lists are then used as targets to train different diagnosis prediction models.
@@ 2 ICD-Related Tasks
We explore two complementary and interconnected tasks: ICD coding, which involves normalizing a physician’s diagnostic conclusion into an appropriate set of ICD codes, and diagnosis prediction, where we aim to identify multiple diagnoses based on a patient’s complete medical history without relying on doctor-assigned diagnoses (Fig. 2).
@@
is akin to medical concept normalization (or entity linking), where the objective is to assign a set of unique ICD codes to the latest patient appointment based on textual diagnosis conclusion written by a doctor. The task is aimed at helping a doctor normalize diagnosed diseases to a set of codes from the complex formal ICD hierarchy. We formulate the ICD Coding task as a modification of traditional information extraction pipeline with three components: (1) Nested Entity Recognition (NER) and (2) Entity Linking (EL) followed by (3) EHR-level code aggregation. Step (3) aims to alleviate the influence of NER component on the resulting quality metrics in an NER+EL pipeline by omitting NER spans. The approach aligns with real-world ICD applications, where the primary objective is accurate assignment of ICD codes (i.e., disease recognition), and imprecise NER outputs are not impactful.
@@
Given an EHR, we perform regular EL over NER predictions. Let $L_{p}=(c^{p}_{1},c^{p}_{2},\dots,c^{p}_{n})$ and $L_{t}=(c^{t}_{1},c^{t}_{2},\dots,c^{t}_{m})$ denote the lists of predicted and ground truth ICD codes, respectively. In standard EL, each list may contain multiple mentions of the same disease (i.e., $c^{t}_{i}=c^{t}_{j}$ for $i\neq j$). Next, we remove duplicates from both lists resulting in unique code sets $S_{p}$ and $S_{t}$ such that $c^{p}_{i}\neq c^{p}_{j}$ and $c^{t}_{i}\neq c^{t}_{j}$ for $\forall_{i\neq j}$. Finally, micro-averaged classification metrics can be computed using the intersection of $S_{p}$ and $S_{t}$ being True Positives ($TP$), and sets differences being False Positives ($FP$) and False Negatives ($FN$) cases: $TP=S_{p}\cap S_{t}$; $FP=S_{t}\setminus S_{p}$; $FN=S_{t}\setminus S_{p}$.
@@
Task: Diagnosis Prediction
is a multi-label classification task also known as automated ICD prediction that outputs likely diagnoses (ICD codes) from a patient’s past medical history, including complaints, test and examination results from previous appointments. In our study, each EHR contains a doctor’s diagnosis conclusion. A major challenge for ICD-grounded applications is that this conclusion is a free-form text, and its normalization to ICD might introduce sensitive errors. Conversely, automatic Diagnosis Prediction is constrained to output ICD-compliant diagnoses by task design.
ICD Coding vs. Diagnosis Prediction
3 ICD Datasets
3.1 RuCCoD: ICD Coding Dataset
@@
Original Dataset Linked Dataset Manual Test Set
@@
The labeling team consisted of three highly qualified experts with advanced education in different fields of medicine, two of whom hold Ph.D. degrees, with every annotation further validated by a fourth expert, a Ph.D. holder in medicine. Grounded in the ICD-10 CM (Clinical Modification) system, the team aimed to identify all nosological units in a diagnosis conclusion and assign the most accurate ICD code to each. An annotation example is shown in Fig. 1. The dataset was randomly split into 3,000 training and 500 testing records. Each expert independently annotated 1,000 training records for diverse labeling, while all three annotated the same 500 test records for consistency. An ICD code was accepted if at least two annotators agreed. Annotation guidelines are in Appx. A.
Inter-Annotator Agreement
We assessed annotation consistency among experts using the Inter-Annotator Agreement (IAA) metric, defined as the ratio of accepted codes to the total unique codes assigned per record Luo et al. (2019). Among ICD codes, the IAA value was 50%, indicating moderate agreement. We note that the inherent subjectivity in manual ICD coding is evident in the low inter-coder agreement among human experts. Existing studies have reported fair to moderate agreement on terminal ICD codes, with Kappa values ranging from 27% to 42%, corresponding to agreement rates of 29.2% and 46.8%, respectively Stausberg et al. (2008). The reported accuracy of coding exhibits significant variability, ranging from 53% to 98% Campbell et al. (2001) and from 41.8% to 88.87% Hosseini et al. (2021). To sum up, the evaluation of inter-annotator agreement using the IAA metric highlights both the challenges and strengths of our annotation process.
@@ 3.2 RuCCoD-DP: Diagnosis Prediction Dataset
RuCCoD-DP includes doctor appointments from 2017 to 2021, divided into four parts: (i) patient complaints and anamnesis, (ii) lab test results, (iii) appointment summary (including assigned ICD codes), and (iv) appointment history. Although RuCCoD and RuCCoD-DP share a common source, we ensure both sets to have no overlapping appointments and patients.
@@
ICD has a fine-grained disease hierarchy introducing a significant challenge even for a qualified doctor to formalize a correct general diagnosis to a specific ICD code. For instance, a H10 Conjunctivitis disease group has 8 specifications including: H10.0 mucopurulent, H10.1 acute atopic, H10.2 other acute, and H10.3 unspecified acute conjunctivitis. Thus, doctor-assigned ICD codes in real-world EHRs can expose substantial errors even if a disease is diagnosed correctly. Our work is the first to explore the mismatch between verbose diagnosis conclusion and the assigned ICD codes by modeling a virtual AI-based medical expert. For each EHR, we consider two ICD code sets: (i) real-world ICD codes originally written by physicians within the EHR (doctor-assigned codes); (ii) automatically assigned ICD codes predicted by a neural model trained on RuCCoD (AI codes). To obtain AI codes, we applied an automated ICD coding for each appointment to relabel the ICD codes based on a doctor’s appointment summary. The AI-generated codes capture medical entities explicitly mentioned in an appointment report.
@@
The collection of two sets of labels allows exploration of whether manual or generated ICD labels are more reliable for model training. For a fair comparison of the labeling approaches, we manually labeled a common test set from a subset of the original appointment dataset’s test set. We formed it by selecting a subset from the test part of the original appointment dataset. For annotation, we adopted the same annotation methodology as for the RuCCoD dataset (Sec. 3.1). The IAA between the doctors was 50% for exact ICD codes and 74% for ICD groups. The final statistics for original, linked datasets as well as the common manual test is summarized in Tab. 2.
@@ 4 ICD Coding Evaluation
For ICD coding experiments, we experiment with the following approaches: 1) a fine-tuned BERT-based pipeline for information extraction, 2) a large language model (LLM) with Parameter-Efficient Fine-Tuning (PEFT), and 3) LLM with retrieval-augmented generation (RAG). All three systems use the same dictionary, with 17,762 pairs of codes and diagnoses (refered to as ICD dict). This was compiled using data from the Ministry of Health. In addition, LLM-based systems used a train set as a dictionary as well. See the Appx. G for a list of the LLMs used. See related work in Appx. B.
@@ 4.1 Models
We developed our information extraction (IE) system with two sequential modules for each task. The NER module with the softmax layer extracts relevant entities, while the EL module links these extracted entities to corresponding ICD codes. For NER, we utilize the pre-trained RuBioBERT Yalunin et al. (2022), and for EL, we employ the multilingual state-of-the-art models SapBERT (Liu et al., 2021a, b), CODER (Yuan et al., 2022b), and BERGAMOT Sakhovskiy et al. (2024). We fine-tuned models on training EL sets via synonym marginalization as suggested by the authors of BioSyn Sung et al. (2020). For more details, see Appx G.
@@
We explored the capabilities of LLMs for clinical coding using PEFT with Low-Rank Adaptation (LoRA) Hu et al. (2021). The pipeline comprised two stages: NER and EL, following the same structure as the BERT-based IE pipeline described earlier. The NER stage involved fine-tuning models on RuCCoD using task-specific prompts (Appx. H). Predictions on the test set were validated via exact string matching or Levenshtein distance (threshold $\leq 2$) to accommodate spelling variations. For the EL stage, a RAG approach was employed to link extracted entities to ICD codes. This approach involved three strategies for building the retrieval component: (i) Using only the ICD dict with embeddings from the widely recognized BGE model Chen et al. (2024), (ii) Combining the ICD dict and RuCCoD training entities, also with BGE embeddings, and (iii) Fine-tuning BERGAMOT embeddings Sakhovskiy et al. (2024) on RuCCoD while using the ICD dict.
The retrieval process used a FAISS index Douze et al. (2024) for each strategy. For each entity extracted in the NER stage, the top-15 most similar entries were retrieved from the index. The final ICD code was assigned using an LLM to select the closest match from the retrieved candidates (prompt in Appx. H). To address class imbalance, diagnosis lists were shuffled during training, forcing models to learn contextual code-discrimination. Fine-tuning parameters followed standard LoRA configurations (Tab. 4, Appx. G).
@@
As an ablation study, we evaluated the same pipeline as in the PEFT stage but without fine-tuning to isolate the LLMs’ inherent capabilities. We used only the fine-tuned BERGAMOT embeddings from strategy (iii) for retrieval, retaining the FAISS index and prompts (Appx. H). The LLM selected ICD codes from retrieved candidates if no direct match was found, replicating the EL process from the PEFT stage. This setup allowed us to quantify the contribution of fine-tuning versus zero-shot inference.

### v2

@@ Abstract
This study investigates the feasibility of automating clinical coding in Russian, a language with limited biomedical resources. We present a new dataset for ICD coding, which includes diagnosis fields from electronic health records (EHRs) annotated with over 10,000 entities and more than 1,500 unique ICD codes. This dataset serves as a benchmark for several state-of-the-art models, including BERT, LLaMA with LoRA, and RAG, with additional experiments examining transfer learning across domains (from PubMed abstracts to medical diagnosis) and terminologies (from UMLS concepts to ICD codes). We then apply the best-performing model to label an in-house EHR dataset containing patient histories from 2017 to 2021. Our experiments, conducted on a carefully curated test set, demonstrate that training with the automated predicted codes leads to a significant improvement in accuracy compared to manually annotated data from physicians. We believe our findings offer valuable insights into the potential for automating clinical coding in resource-limited languages like Russian, which could enhance clinical efficiency and data accuracy in these contexts. Our code and dataset are available at https://github.com/auto-icd-coding/ruccod.
@@
Aleksandr Nesterov1thanks: These authors contributed equally to this work., Andrey Sakhovskiy2,3, Ivan Sviridov4, Airat Valiev5 Vladimir Makharev1,6, Petr Anokhin1, Galina Zubkova4, Elena Tutubalina1,2,7 1 AIRI 2 Sber AI 3 Skoltech 4 Sber AI Lab 5 HSE University 6 Innopolis University 7 ISP RAS Research Center for Trusted AI
@@ 1 Introduction
Although ICD code assignment is crucial for EHRs, it poses significant challenges. Human coders must navigate a wide array of medical terminology, subjective interpretations, and time pressures, all while staying updated with constantly changing classification standards Burns et al. (2012); O’Malley et al. (2005); Cheng et al. (2009). Coding errors can lead to misdiagnosis, ineffective treatment, diminished trust in the healthcare system, and negative public health outcomes. Furthermore, errors in manual coding in the ICD system, result in financial repercussions, accounting for 6.8% of the total payments Manchikanti (2002).
@@
Figure 2: An overview of ICD coding and diagnosis prediction tasks. Diagnosis prediction (left, yellow) is to predict diagnoses as ICD codes based on prior EHR data and current visit details, excluding a doctor’s conclusion. Two training label types are compared: (i) original ICD codes (manually assigned by physicians, shown in green) and (ii) AI ICD codes (automatically generated by our ICD coding pipeline, shown in blue). ICD coding pipeline (right) extracts and links diseases from a doctor-written conclusion, then assigns the deduplicated codes to an EHR.
Schematic description of the ICD coding and diagnosis prediction tasks (left panel, blue and yellow, respectively), along with a detailed illustration of the ICD coding pipeline (right panel).
@@
Despite extensive research on ICD coding using neural networks Li and Yu (2020); Zhou et al. (2021); Yuan et al. (2022a); Baksi et al. (2024); Boyle et al. (2023); Mullenbach et al. (2018a); Cao et al. (2020); Yuan et al. (2022a); Yang et al. (2022); Huang et al. (2022), significant challenges persist for non-English languages. These include low inter-coder agreement, limited labeled data, variability in clinical notes, the hierarchy of ICD codes, and reliance on incomplete input data. In addition, clinical environments in low-resource languages exhibit a shortage of certified medical coders, forcing physicians to assign ICD codes themselves, leading to inconsistencies and low-quality coding. To address these issues, we introduce a novel dataset for automatic ICD coding in Russian and explore automated approaches to assist physicians in the coding process.
@@
Previous studies have primarily focused on English-language datasets, specifically MIMIC-III/IV Goldberger et al. (2000); Johnson et al. (2023). Despite being one of the top ten languages in terms of concept name count within the Unified Medical Language System (UMLS) Bodenreider (2004) biomedical metathesaurus, Russian remains underdeveloped in the clinical domain. The Russian segment of the UMLS comprises only 1.96% of the vocabulary and 1.62% of the source counts found in the English UMLS NIH . Recent corpora, such as RuCCoN Nesterov et al. (2022) and NEREL-BIO Loukachevitch et al. (2023), focus on concepts within the Russian UMLS.
In this work, we explore two closely related tasks: ICD coding and Diagnosis Prediction (DP). Both tasks are designed to assist physicians by standardizing and automating the diagnosis process into ICD codes, especially in low-resource clinical settings where professional coders are not available. As seen in Fig. 2, the tasks take non-overlapping input and complement each other: ICD coding normalizes a free-form doctor’s diagnosis conclusion into a set of relevant ICD codes while the DP task is to directly predict ICD-agreed diagnoses from EHRs in one pass without relying on the doctor’s textual diagnosis conclusion. Although we formulate ICD coding as an entity normalization task and DP as multilabel classification, both tasks are sometimes referred to as ICD coding. Unlike prior classification-based ICD coding research (Li and Yu, 2020; Vu et al., 2020; Wang et al., 2024), we explore a more challenging scenario in which a diagnostic model, acting as an independent medical expert, predicts diagnoses from patient data only without relying on the doctor’s diagnosis conclusion. Thus, we term the classification task diagnosis prediction as it better reflect the problem’s nature and does not create a confusion with linking-based ICD coding Lavergne et al. (2016); Névéol et al. (2017); Coutinho and Martins (2022).
@@
For ICD coding, we present RuCCoD (Russian ICD Coding Dataset), a novel dataset in Russian, labeled by medical professionals based on concepts from the ICD-10 CM (Clinical Modification) system (Sec. 3.1). Second, we establish a comprehensive benchmark for state-of-the-art models, including a BERT-based Devlin et al. (2019) pipeline for information extraction, a LLaMa-based Touvron et al. (2023) model with Parameter Efficient Fine-Tuning (PEFT) and with retrieval-augmented generation (RAG). Furthermore, we evaluate transfer learning of models trained on UMLS concepts and similar biomedical datasets (PubMed abstracts Loukachevitch et al. (2023), clinical notes Nesterov et al. (2022). The results suggest that the ICD’s fine-grained hierarchical structure hinders generalization from other clinical sources (Sec. 4).
@@
For diagnosis prediction, we perform a set of experiments on RuCCoD-DP, a large in-house dataset of 865k EHRs from 164k patients. When training a diagnostic model, we experiment with ICD codes assigned by doctors during patient appointments, as well as the AI-assigned ICD codes (Sec. 5), that is, diagnoses assigned by automatically linking an EHR diagnosis conclusion with a top-performing ICD coding model on RuCCoD (see Fig. 2). Our experiments have revealed that pre-training on automatically assigned ICD codes gives a huge weighted F1-score growth of 28% for diagnosis prediction compared to physician-assigned ICD codes indicating the difficulty of ICD-guided diagnosis formalization for physicians and great potential of AI-aided diagnosing. Our work provides a foundation and guidance for ICD-related research in low-resource clinical languages.
@@ 2 ICD-Related Tasks
is akin to Entity Linking (EL), where the objective is to assign a set of unique ICD codes to the latest patient appointment based on textual diagnosis conclusion written by a doctor. The task aims to help a physician normalize diagnosed diseases to a set of codes from the complex formal ICD hierarchy. We model the ICD Coding as an information extraction pipeline with three components: (1) Nested Entity Recognition (NER) and (2) EL followed by (3) EHR-level code aggregation. Step (3) minimizes NER influence on pipeline metrics by omitting NER spans. The approach aligns with real-world ICD applications, where the primary objective is accurate assignment of ICD codes (i.e., disease recognition), and imprecise NER outputs are not impactful.
@@
Given an EHR, we perform EL on NER predictions. Let $L_{p}=(c^{p}_{1},c^{p}_{2},\dots,c^{p}_{n})$ and $L_{t}=(c^{t}_{1},c^{t}_{2},\dots,c^{t}_{m})$ denote the lists of predicted and ground truth ICD codes, respectively. After a standard NER+EL pipeline, each list may contain multiple mentions of the same disease (i.e., $c^{t}_{i}=c^{t}_{j}$ for $i\neq j$). The presence of duplicate disease mentions prevents entity-level metric aggregation, as they are erroneously counted multiple times, thereby introducing bias into the performance evaluation. We remove duplicates from both lists, resulting in unique code sets $S_{p}$ and $S_{t}$ such that $c^{p}_{i}\neq c^{p}_{j}$ and $c^{t}_{i}\neq c^{t}_{j},\forall i\neq j$. Finally, micro-averaged classification metrics are calculated from True Positives ($TP$), False Positives ($FP$) and False Negatives ($FN$): $TP=S_{p}\cap S_{t}$; $FP=S_{p}\setminus S_{t}$; $FN=S_{t}\setminus S_{p}$.
Diagnosis Prediction
is a multi-label classification task that outputs likely diagnoses (ICD codes) for the current doctor appointment from a patient’s past medical history, including complaints, test and examination results from previous appointments. In our study, each EHR contains a doctor’s diagnosis conclusion. A major challenge for ICD-grounded applications is that this conclusion is a free-form text, and its normalization to ICD might introduce sensitive errors. Conversely, automatic Diagnosis Prediction is constrained to output ICD-compliant diagnoses by task design.
ICD Coding vs. Diagnosis Prediction
3 ICD Datasets
3.1 RuCCoD: ICD Coding Dataset
@@
Original Dataset Linked Dataset Common Test Set
@@
The labeling team consisted of three highly qualified experts with advanced education in different fields of medicine, two of whom hold Ph.D. degrees, with every annotation further validated by a fourth expert, a Ph.D. holder in medicine. Grounded in the ICD-10 CM (Clinical Modification) system, the team aimed to identify all nosological units in a diagnosis conclusion and assign the most accurate ICD code to each. An annotation example is shown in Fig. 1. The dataset was randomly split into 3,000 training and 500 testing records. Each expert independently annotated 1,000 training records for diverse labeling, while all three annotated the same 500 test records for consistency. An ICD code was accepted if at least two annotators agreed. Annotation guidelines and details on inter-annotator agreement are given in Appx. A.
@@ 3.2 RuCCoD-DP: Diagnosis Prediction Dataset
RuCCoD-DP includes doctor appointments from 2017 to 2021, divided into four parts: (i) patient complaints and anamnesis, (ii) lab test results, (iii) appointment summary (including assigned ICD codes), and (iv) appointment history. Although RuCCoD and RuCCoD-DP share a common source, we ensured that both sets contain no overlapping patients and, consequently, no overlapping appointments.
@@
ICD has a fine-grained disease hierarchy introducing a significant challenge even for a qualified doctor to formalize a correctly diagnosed disease . For instance, a H10 Conjunctivitis disease group has 8 specifications including: H10.0 mucopurulent, H10.1 acute atopic, H10.2 other acute, and H10.3 unspecified acute conjunctivitis. Thus, doctor-assigned ICD codes in real-world EHRs can expose substantial errors even if a general disease is diagnosed correctly. To address the issue, we consider two ICD code sets for each EHR: (i) real-world ICD codes originally written by physicians within the EHR (doctor-assigned codes); (ii) automatically assigned ICD codes predicted by a neural model trained on RuCCoD (AI codes). AI codes (i.e., AI-assigned diseases) are assigned to an EHR by applying our top-performing BERT-based NER+EL ICD Linking pipeline (Tab. 3) to the EHR’s doctor’s real-world diagnosis conclusion (see Fig. 2). Our pipeline extracts diseases (NER) and links (EL) them to ICD codes and then the found diseases are assigned to the given EHR labels for ICD Coding. Thus, AI codes are designed to aid in the formalization of the human-written diagnosis to the ICD code set while relying only on the written conclusion of the physician. Notably, the two coding types rely on the same underlying free-form diagnosis conclusions.
@@
The collection of two sets of labels allows exploration of whether manual or generated ICD labels are more reliable for model training. For a fair comparison of the labeling approaches, we developed a common test set from a subset of the original appointment dataset’s test set. We formed it by selecting a subset from the test part of the original appointment dataset. For annotation, we adopted the same annotation methodology as for the RuCCoD dataset (Sec. 3.1). The mean pairwise JSC across annotators was 0.331. The final statistics for original, linked datasets as well as for the common test set is summarized in Tab. 2.
@@ 4 ICD Coding Evaluation
For ICD coding experiments, we experiment with the following approaches: 1) a fine-tuned BERT-based pipeline for information extraction, 2) a large language model (LLM) with Parameter-Efficient Fine-Tuning (PEFT), and 3) LLM with retrieval-augmented generation (RAG). All three systems use the same dictionary, with 17,762 pairs of codes and diagnoses (refered to as ICD dict) compiled from the Ministry of Health data. In addition, LLM-based systems used a train set as a dictionary as well. See the Appx. G for a list of the LLMs used. See related work in Appx. B.
@@ 4.1 Models
Our Information Extraction (IE) pipeline uses sequential NER and EL modules. The NER module, employing a softmax layer, extracts relevant entities, and the EL module then links these entities to ICD codes based on semantic similarity with ICD dictionary entries. For NER, we utilize the pre-trained RuBioBERT Yalunin et al. (2022), and for EL, we employ the multilingual state-of-the-art models SapBERT (Liu et al., 2021a, b), CODER (Yuan et al., 2022b), and BERGAMOT Sakhovskiy et al. (2024). In EL, entities and vocabular disease names are encoded by a BERT-based model with average pooling, and the top-$k$ concepts are selected for each entity based on the nearest Euclidean distance. We fine-tuned models on EL train sets via synonym marginalization proposed in BioSyn Sung et al. (2020). For more details, see Appx G.
@@
We explored the capabilities of LLMs for clinical coding using PEFT with Low-Rank Adaptation (LoRA) Hu et al. (2021). The pipeline included two steps: NER and EL, following the structure of BERT-based IE pipeline described earlier. For NER stage, models were fine-tuned on RuCCoD using task-specific prompts (Appx. H). The predictions were validated by exact string matching and Levenshtein distance with a threshold $\leq 2$ chosen empirically to optimize the robustness of the spelling without overcorrecting semantically distinct entities. For EL, a RAG approach was implemented to link extracted entities to ICD codes. The retrieval component was built using three strategies: (1) BGE embeddings Chen et al. (2024) on the ICD dict, (2) BGE embeddings on the ICD dict combined with RuCCoD training entities, and (3) BERGAMOT embeddings Sakhovskiy et al. (2024) fine-tuned on RuCCoD with the ICD dict.
We adopted the FAISS index Douze et al. (2024) to retrieve the top-15 most similar dictionary entries for each entity extracted in the NER stage. The final ICD code was assigned using an LLM to select the closest match from the retrieved candidates (prompt in Appx. H). To address class imbalance, diagnosis lists were shuffled during training, forcing models to learn contextual code-discrimination. Fine-tuning parameters followed standard LoRA configurations (Tab. 5, Appx. G).
@@
As an ablation study, we evaluated the same pipeline as in the PEFT stage but without fine-tuning to isolate the LLMs’ inherent capabilities. We used only the fine-tuned BERGAMOT embeddings from strategy (3) for retrieval, retaining the FAISS index and prompts (Appx. H). The LLM selected ICD codes from retrieved candidates if no direct match was found, replicating the EL process from the PEFT stage. This setup allowed us to quantify the contribution of fine-tuning versus zero-shot inference.